OpenStudy (anonymous):

If you were to solve the following system by substitution, what would be the best variable to solve for and from what equation? 3x+6y=9 2x-10y=13 a)x in first equation b)x in second equation << i think its this just making sure c)y in first equation d)y in second equation

hartnn (hartnn):

why you think so ?

OpenStudy (anonymous):

well it is about which would be the best variable to me using 2x-10=13 would be easier

hartnn (hartnn):

you know you can first simplify 3x+6y=9 and then you would try to find which one will be easier

OpenStudy (anonymous):

D: okay ill try that :)

OpenStudy (anonymous):

okay so when simplified it become x+2y=3?

hartnn (hartnn):

right, now which ones easy to isolate x? x+2y = 3 or 2x-10y =13 ?

OpenStudy (anonymous):

the first ....

hartnn (hartnn):

now you got it :)

OpenStudy (anonymous):

:D thanks

hartnn (hartnn):

welcome ^_^